# Twists, Humps, and Pebbles: The Uneven Landscape of End-to-end Speech Recognition and Demographics

### v1

@@ 4 Gender Performance Gaps
Table 1: Model error rate gap (Eq. 1). Positive values indicate better performance on men, negative ones on “other.” $p<0.05$
Restricting our analysis to validation and test sets leaves us with few records for speakers who do not identify as neither male nor female. As such, we report results for five high-resource languages, Catalan, German, English, Spanish, and French. [^12] Size varies from 16 (ca min.) to 86 (es, max) records
Table 1 reports $E(r_{O},r_{M})$ for both models. Results suggest that models penalize non-male speech and that gaps are larger than disparities observed for the female/male groups. Given the limited sample size, we deem this analysis only exploratory. However, it is a valuable first overview of the model’s behaviors for gender-non-conforming voices. We underscore the need for more representation of diverse voices and sociodemographic groups in current resources, and for which monitoring fairness in existing spoken language models remains out of reach. See §7 for an expanded discussion.
5 Acoustic Analysis
Figure 3: Acoustic features. Speaking rate, pitch average and intensity average (left to right). Italian (2(a)) and Slovak (2(b)) on our three datasets. “Male” (green), “Female” (red), and “Other” (blue, available for CV).
Our findings (§4) reveal performance gaps across gender groups, with variable trends over languages and datasets. To further investigate potential reasons, we explore whether phonetic features, as measured across different groups in our evaluation records, shed light on the phenomenon (Q2). For each segment, we examine three parameters: i) pitch average, the mean of the fundamental frequency values measured in Hertz (Hz) Hirst and de Looze (2021); ii) intensity average, the mean of the fundamental frequency values measured in decibel (dB) for each sample Pausewang Gelfer and Young (1997); iii) speaking rate, computed as the number of tokens per minute Künzel (2013). [^13] We compute average pitch and intensity using Praat Boersma and Weenink (2001). Concerning speaking rate, we found no reliable multilingual tools for syllable segmentation. Therefore, we approximated it by the number of tokens as computed by Whisper’s tokenizer.
We analyze these features because they represent sociophonetic variability of interest for the gender dimension and within our setting.
As already discussed in §2, differences in pitch between men and women are well documented Coleman (1976); Hillenbrand and Clark (2009), and variations in speaking rate have also been associated with gendered speech Borsel and Maesschalck (2008). Besides, the selected features might capture distinctive acoustic aspects across the datasets used in our experiments. For instance, intensity can be influenced by the recording conditions Maryn and Zarowski (2015), which can vary significantly in crowdsourced speech data. Additionally, speech rate tends to be higher in spontaneous speech Nakamura et al. (2008). Differences in acoustic parameters may also exist across languages. For example, languages can vary in the average speaking rate Coupé et al. (2019) and pitch Loveday (1981); Yuasa (2008).
@@ 5 Acoustic Analysis
Since we found similar distributions across languages – and due to space constrains – in Figure 3 we report results for two representative languages only: Italian and Slovak. The former exemplifies the set of languages with comparable performance across genders, whereas the latter is biased towards men and belongs to the set of languages with larger gender gaps trends (§4).
By comparing these two groups of languages, pitch emerges at the only feature that differ across female and male, whereas no substantial variation is attested for intensity and speaking rate. Interestingly, the “other” group is slightly more spread on pitch and intensity compared to “male” and “female”. By comparing datasets, we still find similar patterns, except for the intensity parameter. Intensity values are more divergent in their distributions, which is particularly noticeable for Fleurs. [^14] This could be attributed to the underlying recording conditions of Fluers, details of which are however not provided.
Overall, however, we do not attest to significant acoustic differences across datasets and languages, regardless of the gap size or direction (male, female, other). Thus, these findings suggest that acoustic features cannot explain gender-based performance gaps and their trends.
@@ 6 Probing Gender in ASR Models
Since models exhibit gender-based disparities and acoustic features do not fully explain the phenomenon, we turn to interpretability analysis to understand whether models encode records differently across genders.
@@
Motivated by this line of research, we measure gender extractability in ASR models and ask (Q2) whether and to which extent it explains gender-based performance gaps. Ours is the first study of this kind. [^15] Interpretability for speech models is a relatively new research avenue. Mohebbi et al. (2023) use context-mixing techniques to explain homophone disambiguation; Pastor et al. (2023) use word-segment masking to explain semantic classication tasks.
@@
We focus on one dataset-model configuration, namely Whisper and CV. We attach our probes to the model encoder’s last layer embeddings and train one distinct probe for every position. We use Logistic Regression and Minimum Description Length (MDL) (Voita and Titov, 2020) to probe gender in the female-male binary setup (see Appendix A.6 for details).
Figure 4: F1 Macro (top) and measured code length ($1e^{6}$, bottom) for Logistic Regression and MDL probes, respectively. English, F-M setup, first 10 seconds of context. Actual score (pale line), exponential moving average (solid), and average length of test snippets (vertical dashed).
@@
Figure 4 shows gender extractability on the English CV split for the F-M setup. Trends are similar across languages (full results in Figure 8(a)). First, F1 scores suggest that gender extractability is relatively easy. Whisper produces representations of female and male speech that can be easily separated. Second, extractability is not constant over time. It starts high in the first milliseconds, drops during the actual signal (i.e., the speaker is talking), and finally plateaus around the initial value. This finding recalls attention “sink” theory Xiao et al. (2023) implicating that transformers prioritize initial positions for specific tasks. Third, MDL probes share similar trends—low codelength indicates an “easier” task / higher extractability—suggesting that simple logistic regression is a robust method to start with, countering issues found instead in NLP (Hewitt and Liang, 2019; Voita and Titov, 2020). [^16] To confirm this hypothesis, we conduct additional experiments by training logistic probes with randomly shifted labels. We report random guess performance on all languages (see Figure 8(b)) in Appendix A.6.
@@
[…]
@@ 7 Discussion
Our extensive analysis confirms that multi-task multilingual ASR models display gender performance gaps. Depending on the language, disparities can favor male or female speakers, but rarely those who do not identify as either. In the following, we report two discussion points to help frame our study, its findings, and its implications.
Figure 6: Whisper distribution of error rates. Female and male records in Fleurs. nl (top) and ar (bottom). Proportion of records (y axis) having a WER error rate (x axis) at least that large.
Interpreting $E(\cdot)$.
For our gender gap analysis, we study performance disparities by studying the relative difference ($E(\cdot)$) in scores across gender groups (see §3.1). While such definition allows to compare models and languages at different absolute error levels, relative differences might tell little about the concrete impact of gender disparities in ASR. [^18] For instance, a 10% F-M relative difference corresponds to 0.2 or 2 WER points gap if we consider a model that scores 2 or 20 absolute WER on men, respectively. Ideally, a more informative index would indicate the usability of an ASR system, thus providing a more precise estimation of its failures and their severity.
Sentence-level error metrics can provide such a fine-grained view. Following Koenecke et al. (2020), in Figure 6, we report the proportion of generated transcriptions whose quality is greater than given WER values. Due to space constraints, we only show the results for Whisper on Fleurs for two languages as exemplars (i.e., Arabic and Dutch). This visualization lets us identify where performance between genders starts to diverge. Let us hypothesize that transcriptions became unusable after WER>20. The plot can help us spot that men and women would be equally affected in Dutch. However, 50% of Arabic transcripts from men would be unsuitable compared to 40% of women’s. Parity would be achieved if we lose the constraint up until around WER=40.
We echo the literature that urges the need for more informative/concrete bias measurement (Delobelle et al., 2022; Orgad and Belinkov, 2022) and advocate measures that can be used to guide model design and inform its reliable deployment. Thus, this evaluation is particularly relevant and a more realistic estimate of the impact of sociodemographic disparities in real-world settings.
Group-Aware Dataset Splits.
Numerical and statistical studies like ours heavily depend on quality and quantity of evaluation data. In this specific case, we would want to have representative splits, i.e., records from many speakers from each group. However, current dataset releases are drastically skewed. For instance, in most of our datasets and language configurations, men are overrepresented in the number of records and speakers. While downsampling can partially solve the former at evaluation time, little can be done with the latter. Moreover, we point out that stratification over gender and speakers to create splits (train/validation/test) affects languages differently. In a high-resource language, an 80/10/10 division still contributes many speakers from all groups in validation and testing. However, in a low-resource setup, only a few speakers are included—in CV most languages counted only one to 20 speakers considering men and women, and frequently none or less than five speakers for “other”.
We argue that dataset releases should factor in the difficulty in collecting samples from less represented groups and provide group-aware splits to facilitate fairness research.
@@ 8 Conclusion
We conducted the first extensive evaluation on gender-based performance gaps of Whisper (Radford et al., 2022) and SeamlessM4T (Communication et al., 2023) for ASR. These models consistently exhibit gender bias across 19 languages from seven language families. Depending on the language, disparities can favor men or women but rarely favor speakers who identify as neither. We locate a potential source of these gaps using gender probes from interpretability approaches. Our results show that probes can be a proxy for gender gaps and that group fairness in multi-task multilingual models remains unsolved.
@@
The use of gender as a variable in this warrants some ethical reflection.
@@ Methods.
Since we miss reliable multilingual phonetic tools, we estimated the speaking rate using Whisper’s pre-trained tokenizer. We acknowledge that pretrained tokenizers yield fewer tokens for high-resource languages. Hence, our measurement of low-resource might be overestimating the phenomenon.
@@ Appendix A Experimental Details
(c) Common Voice
Figure 7: Seamless and Whisper transcription quality. WER results are computed on test splits.
@@ A.1 Multilingual ASR Models
Table 2: WER across datasets splits (test/train/validation), all presampled female/male records, and on the sampled gender subset used in our experiments. Results averaged over 19 languages in CV and Fleurs, and 11 for VoxPopuli.
@@
Table 3: Average F-M WER absolute difference across datasets splits (test/train/validation), all presampled female/male records, and on the sampled gender subset used in our experiments. The results are averaged over 19 languages for CV and 11 for VoxPopuli. For Fleurs, we average results over the only 9 languages that comprise both female/male speakers in all splits.
A.2 Voice Activity Detection
@@
A.3 Speaker ID Attribution in Fleurs
Figure 8: UMAP projections of snippets from men (left, 13 speakers found) and women (right, 16 speakers found) in Fleurs-en. Color indicates cluster ID assigned by HDBSCAN.
@@ A.3 Speaker ID Attribution in Fleurs
A.4 Subset selection and data splits
To ensure the soundness of our experimental framework and of the gender subsample selection described in in Section 3, we conducted preliminary analyses by comparing WER transcription quality scores (Table 2) and gender F-M gap results (Figure 3) across: i) different datasets splits (test/train/validation), ii) all female/male records available from any split, and iii) the sampled female/male subset used in our main experiments.
@@ A.4 Subset selection and data splits
Figure 2 shows no notable WER variations for Fleurs. Instead, both models achieve better performance on the train splits of CV and VP, especially for the former dataset. The influence of the split accounted in CV and VP can thus be reflective of the F-M WER differences reported in Figure 3. Informed by such results, the gender subset used in our experiments is sampled from the test and validation split only for CV and VP, whereas for Fleurs we leverage gender data from the whole corpus.
A.5 Dataset Statistics
Lang # records Seconds # Tokens # M Gini (M) # F Gini (F)

### v2

@@ 5.1 Acoustic Analysis
Table 1: Model error rate gap (Eq. 1). Positive values indicate better performance on men, negative ones on “Other.” $p<0.05$.
Motivated by sociophonetic evidence for gendered differences in speech (§2), we examine voice and language production across speakers of different genders. Specifically, we measure three acoustic features in our evaluation records, and explore whether potential acoustic differences across gender groups relate to gender performance gaps.
We include: i) pitch, measured as the mean of the fundamental frequency values Hirst and de Looze (2021), known to vary between biological sexes Coleman (1976); Hillenbrand and Clark (2009) and languages (Loveday, 1981; Yuasa, 2008); ii) intensity, measured as the mean of fundamental frequencies Pausewang Gelfer and Young (1997), subject to recording conditions Maryn and Zarowski (2015); iii) speaking rate, measured as the number of tokens per minute Künzel (2013), subject to language and gender variation (Borsel and Maesschalck, 2008; Coupé et al., 2019) as well as read vs. spontaneous speech Nakamura et al. (2008). [^11] We compute average pitch and intensity using Praat Boersma and Weenink (2001). Concerning speaking rate, we found no reliable multilingual tools for syllable segmentation. Therefore, we approximated it by the number of tokens as computed by Whisper’s tokenizer.
@@
We started by exploring whether pitch, intensity, and speaking rate differ between genders in each dataset-language pair. Independent-sample T-tests revealed significant differences for most setups ($p<0.05$), consistent with previous literature. For example, pitch consistently showed statistically significant differences (see Appendix B.1).
Motivated by these observations, we computed the mean acoustic values of each gender group and correlated it with performance gaps, i.e., $E(r_{F},r_{M})$, aggregating by dataset and model. No clear trends emerged overall. Sporadically, we found a strong and significant linear correlation, e.g., for Seamless and VoxPopuli where Pearson’s rho for pitch is $-0.83$ (see Appendix B.1 for complete details). Fitting an ordinary least squares (OLS) regressor to predict $E(r_{F},r_{M})$ led to similar conclusions, as low R2 scores suggest (max: $0.42$, averageσ: $0.24_{\pm 0.11}$). For a more fine-grained analysis, we repeated the analysis fitting an OLS model to predict sentence-level error rates (i.e., $r_{F}$, $r_{M}$) but found low R2 scores (max: $0.20$, averageσ: $0.03_{\pm 0.04}$).
In summary, despite pitch, intensity, and speaking rate vary significantly across gender groups within the same dataset and language, such variation seldom correlates with performance differences. These findings suggest that performance gaps are a complex phenomenon, and our analysis should expand elsewhere to explain it.
5.2 Lexical Analysis
In §5.1, we explored acoustic aspects of speech. However, what is uttered can be as crucial as how it is uttered. Indeed, prior work has found that overall speech perception can be impacted by aspects related to lexical and syntactic complexities (van Knijff et al., 2018; Carroll and Ruigendijk, 2012). Besides, certain lexical phenomena such as named entities represent a well-known challenge for speech models, especially in multilingual contexts (Gaido et al., 2021). Thus, we focus on speech content and study whether lexical phenomena—as measured in the reference transcripts—explain ASR disparities.
For each record, we counted i) the occurrences of part-of-speech tags and ii) named entities, and computed iii) lexical density (Halliday, 1989). Similarly to acoustic features, we contrast distribution between gender groups (details in Appendix B.2).
Findings.
Comparing the distribution of lexical features between female and male speakers, we found mixed results. Several dataset-language setups have significantly different distributions ($p<0.05$), but not all. Different distributions are also present in read datasets, suggesting that lexical variability is not controlled when collecting data. We proceeded similarly to the acoustic analysis to verify whether such differences explain error rates. Fitting an OLS model to predict error rates from lexical features only yields low R2 scores (max: 0.35, averageσ: 0.09±0.08, across all datasets and languages). Moreover, we found no significant (linear) correlation between the difference of group means and $E(r_{F},r_{M})$.
This finding echoes those from acoustic analysis: Lexical phenomena do not explain gender performance gaps in our data. We must research other aspects beyond the lexical content of utterances.
@@ 6 Probing Gender in ASR Models
Motivated by this line of research, we measure gender extractability in ASR models and ask (Q2) whether and to what extent it explains gender-based performance gaps. To our knowledge, ours is the first study of this kind. [^12] Interpretability for speech models is a relatively new research avenue. Mohebbi et al. (2023) use context-mixing techniques to explain homophone disambiguation; Pastor et al. (2023) mask word-units to explain intent detection models.
@@
We focus on one dataset-model configuration, namely Whisper and CV. We attach our probes to the model encoder’s last layer embeddings and train one distinct probe for every position. We use Logistic Regression and Minimum Description Length (MDL) (Voita and Titov, 2020) to probe gender in the female-male binary setup (see Appendix B.3 for details).
Figure 3: F1 Macro (top) and measured code length (bottom) for Logistic Regression and MDL probes, respectively. English, F-M setup, first 10 seconds of context. Actual score (blue pale line), exponential moving average (solid line, n=3), and average length of test snippets (dashed line).
@@
Figure 3 shows gender extractability on the English CV split for the F-M setup. Trends are similar across languages (full results in Figure 9(a)). First, F1 scores suggest that gender extractability is relatively easy. Whisper produces representations of female and male speech that can be easily separated. Second, extractability is not constant over time. It starts high in the first milliseconds, drops during the actual signal (i.e., the speaker is talking), and finally plateaus around the initial value. This finding recalls the “attention sink” theory, implicating that transformers prioritize initial positions for specific tasks Xiao et al. (2023). Third, MDL probes share similar trends—low codelength indicates an easier task / higher extractability—suggest that simple logistic regression is a robust method and countering the issues found in NLP (Voita and Titov, 2020). [^13] To confirm this hypothesis, we conduct additional experiments by training logistic probes with randomly shifted labels. We report random guess performance on all languages (see Figure 9(b)) in Appendix B.3.
@@
Do gender probing scores tell us something about ASR quality and gender bias (Q2)? To answer this question, we compare F1 scores from logistic regression scores and ASR error rates. We measure F1S, i.e., the average F1 score achieved by probes at the positions corresponding to actual speech, [^14] We approximate this interval by considering all positions before the average test set length (orange line in Figure 3). $\phi(r_{M})$ and $\phi(r_{F})$ (error rates on male and female segments), and the disparity score $E(r_{F},r_{M})$.
F1S has a weak linear correlation with $\phi(r_{M})$ and $\phi(r_{F})$ (Pearson’s $\rho$ is 0.13 and -0.003, respectively), hinting that F1S cannot explain per-group quality. However, $\rho$ between F1S and $E(r_{F},r_{M})$ is -0.65 ($n=19$, see Figure 8 in Appendix B.3). In other words, linear correlation suggests that the better probes can extract gender from Whisper’s hidden states, the lower the F-M error rate gap is. Note that, as per its definition, $E(r_{F},r_{F})$ can be negative (see points below the zero line in Figure 8). Negative values indicate performance favoring the minority group (here, women). This finding suggests that Whisper does encode recordings from speakers that identify as men and women differently. This aspect can serve as a proxy for measuring and mitigating gender disparities. However, we caution the reader from attributing high extractability to gendered voice and discourage using probes on people’s voice to predict gender (see Ethical considerations, §Ethical Considerations).
@@ 7 Discussion
Twists, humps, and pebbles—the way to equitable multilingual speech recognition models is no straight line. We have discovered that gender disparities vary across spoken setups and languages and that consistency is also weak between models. Perhaps of greater interest, the advantaged group is sometimes women, sometimes men, and never anyone who does not identify with either.
Such a high variability prevents us from foreseeing gender disparities and, in turn, the actual impact on system users. We discuss here how to i) reduce such variability and ii) better use the metrics and test data at our disposal.
Free! the training data.
While our acoustic and lexical inquiries on test data were inconsequential, experiments with probes on the model’s internals showed that models learned and encoded properties to differentiate genders. Intuitively, a follow-up study would focus on what shaped those model’s properties in the first place: training data. Models encode facts and information from training data, and some studies looked into how different sociodemographic groups are represented (Muller et al., 2023; Elazar et al., 2024). Studying gender distributions in pretraining multilingual speech data would be extremely valuable, as we could establish how gendered gaps and underrepresentation correlate. However, this analysis was impossible, as neither Whisper (Radford et al., 2022) nor SeamlessM4T (Communication et al., 2023) released such information. We call for more transparency.
Figure 4: Share of records (y axis, %) having a WER greater than the value on the x axis. Whisper, Fleurs nl (top) and ar (bottom).
Go beyond group-wise metrics.
A crucial choice in our design was measuring differences between two groups of speakers. We opted for $E(\cdot)$, measured as the relative difference between the group means. While the metric was primarily meant to make results across datasets, languages, and models comparable, it does not provide a complete picture of gender disparity’s concrete impact. For instance, it hides absolute gaps: a $E(\cdot)=10$ gap corresponds to a 0.2 WER gap if the model averages a WER of 2 on male speakers and to a 2 WER gap if the average is 20. These gaps may or may not be significant given many contextual factors (e.g., what the model is being used for, who is using it, and which ramifications such gaps can have). This thinking echoes that of recent critiques on measuring social bias in NLP, advocating for explicit statements about what system behavior must be considered good or bad and under which social values (Blodgett et al., 2020), and how to build contextualized metrics (Lum et al., 2024).
We would like a metric to indicate a more precise estimation of ASR model failures and their severity for users. A step in this direction would be going beyond group-wise metrics and studying sentence-level errors. Drawing on the concepts of wealth distribution and inequality (O’sullivan et al., 2003), we can inspect error distributions between speakers and groups and address the question “assuming X is an acceptable error rate, what fraction of records from each group will be served to a satisfactory level?” Following Koenecke et al. (2020), we report in Figure 4 the proportion of records having an error rate at least that large. Due to space constraints, we focus on Whisper’s transcriptions of Fleurs’ Arabic and Dutch data. This visualization lets us identify where performance between genders starts to diverge. If, given our use case, we hypothesize that transcriptions become unusable for WER > 20, men and women would be equally affected in Dutch. However, it would not be the case in Arabic where 40% of records from female speakers, but half of those from males, would not be acceptable. Parity is reached around WER=40.
This concrete analysis highlighted gender disparities that group-wise metrics could not capture otherwise. It also served as an example of a broader argumentation: Future inquiries can benefit from using metrics and observations more closely related to real-world scenarios.
Improve sociodemographic representation.
Empirical studies like ours rely the quality and quantity of evaluation data. To test speech recognition, we need representative splits from diverse speakers. However, current datasets are gender-skewed, with men often overrepresented in terms of records and speakers. While downsampling can help at evaluation time, it does not address the lack of speaker diversity. Skewed distributions are even more evident in low-resource languages when building standard train/validation/test sets is required—e.g., in CV, Yoruba counts only 14 and 20 unique males and females, respectively. Catalan counts only 8 within “Other.”
Therefore, we call for much needed considerations when comparing gender groups in high-resource languages, and new collective efforts to collect more representative data for low-resource languages and gender identities beyond then binary.
@@ 8 Conclusion
We conducted the first extensive evaluation on gender-based performance gaps of Whisper (Radford et al., 2022) and SeamlessM4T (Communication et al., 2023) for ASR. These models consistently exhibit gender bias across 19 languages from eight language families. Depending on the language, disparities can favor men or women but rarely favor speakers who identify as neither. We locate a potential source of these gaps using gender probes from interpretability approaches. Our results show that probes can be a proxy for gender gaps and that group fairness in multi-task multilingual models remains unsolved.
@@
The use of gender as a variable in this paper warrants ethical reflections.
@@ Methods.
Since we miss reliable multilingual tools for acoustic analyses, we estimated the speaking rate using Whisper’s pre-trained tokenizer. We acknowledge that pretrained tokenizers yield fewer tokens for high-resource languages. Hence, our measurement of low-resource might be overestimating the phenomenon.
@@ Appendix A Experimental Details
(c) CV
Figure 5: Seamless and Whisper transcription quality. Error rate results are computed on test splits.
@@ A.1 Multilingual ASR Models
Table 2: Error rate distribution across datasets splits (test/train/validation), all pre-sampled female/male records, and on the sampled gender subset we used. Results averaged over 19 languages in CV and Fleurs, and 11 for VoxPopuli.
@@
Table 3: Average F-M ER absolute difference across datasets splits (test/train/validation), all pre-sampled female/male records, and on the sampled gender subset used in our experiments. The results are averaged over 19 languages for CV and 11 for VoxPopuli. For Fleurs, we average results over the only 9 languages comprising female/male speakers in all splits.
A.2 Languages
Language ISO Family
Japanese ja Japonic
Dutch nl Germanic
English en
German de
Swahili sw NC / Bantu
Yoruba yo NC / Volta-Niger
Catalan ca Romance
French fr
Italian it
Portuguese pt
Romanian ro
Spanish es
Arab ar Semitic
Czech cs Slavic
Russian ru
Serbian sr
Slovak sk
Finnish fi Uralic
Hungarian hu
Table 4: Languages, their ISO code, and family studied in this paper.
We studied 19 languages in CV and Fleurs and 11 in VoxPopuli. Languages cover eight distinct language families. Due to data availability, we limited the comparison between speaker identifying with “Male” and “Other” in CV to Catalan, German, English, Spanish, and French. Table 4 reports an overview of languages and language families.
A.3 Voice Activity Detection
@@
A.4 Speaker ID Attribution in Fleurs
Figure 6: UMAP projections of snippets in Fleurs-en. We found 13 male speakers (left) and 16 female speakers (right). Color indicates cluster ID assigned by HDBSCAN.
@@ A.4 Speaker ID Attribution in Fleurs
A.5 Subset selection and data splits
To ensure the soundness of our experimental framework and of the gender subsample selection described in Section 3, we conducted preliminary analyses by comparing WER transcription quality scores (Table 2) and gender F-M gap results (Figure 3) across: i) different datasets splits (test/train/validation), ii) all female/male records available from any split, and iii) the sampled female/male subset used in our main experiments.
@@ A.5 Subset selection and data splits
Table 2 shows no notable WER variations for Fleurs. Instead, both models perform better on the train splits of CV and VP, especially for the former dataset. The influence of the split accounted in CV and VP can thus be reflective of the F-M WER differences reported in Figure 3. Informed by such results, the gender subset used in our experiments is sampled from the test and validation split only for CV and VP, whereas for Fleurs, we leverage gender data from the whole corpus.
A.6 Dataset Statistics
Lang # records Seconds # Tokens # M Gini (M) # F Gini (F) # O Gini (O)